Two ships leave port at the same time. Ship X is heading due north and Ship Y is heading due east. Eight hours later they are 400 miles apart. If the Ship X had traveled 320 miles from the port, how many miles had Ship Y traveled?
Found 2 solutions by richwmiller, amey1309:
Answer by richwmiller(17219)

one leg is 320
the hypotenuse 400
320+a^2=400^2
a=399.6

Answer by amey1309(2)

NORTH
|\
| \
|X \(400)
| \
|320 \
| \
|______\EAST
Y
(HY)^2=(X)^2+(Y)^2
(400)^2=(380)^2+(Y)^2
Y=((400)^2-(320)^2)^1/2
Y=(57600)^1/2
Y=240
Therefore Ship Y has Travelled 240 miles